Sketch the graph of a parabola and give the vertex, focal point, and directrix.
A) x^2-6x-3y+3=0
B) 3x^2-12x+2y-10=0

Answer by lynnlo(4176)

(A)X^2-6X-3Y+3=0
(3,-2)=======(3,-5/4)===========Y=-11
VERTEX=======FOCI==============DIRECTRIX
(B)3X^2-12X+2Y-10=0
(2,11)=======(2,65/6)==========Y=67/6
VERTEX======FOCI==============DIRECTRIX
IN CASE YOU DON'T KNOW HOW TO GRAPH YOUR PROBLEM
SEND ME AN E-MAIL AND I WILL SEND THE GRAPH TO YOU